OpenStudy (anonymous):

ok i figured out the substitution. how do i do elimination? Solve –x + 2y = 12 using the elimination method. x + 6y = 20

OpenStudy (anonymous):

add both eqns 8y = 32 y = 4 put in x +6y =20 x + 24 = 20 x = -4 x= -4 and y = 4

OpenStudy (angela210793):

u add both parts ---> x-x+2y+6y=12+20 8y=32 y=32/8=4 then u replace y with 4 at 1 of those eq:X+6*4=20 X=-4

OpenStudy (anonymous):

Solve 9x + 2y = 2 by the elimination methods. 21x + 6y = 4 and for that problem would i combine all the signs too" or should i pick one

OpenStudy (angela210793):

9x + 2y = 2 In this case u multiply the 1sr eq with a (-3) and u'll have 21x + 6y = 4 -27x-6y=-6 21x+6y=4 (+) -6x=-2 x=1/3 21*1/3+6y=4 7+6y=4 6y=-3 y=-1/2

OpenStudy (angela210793):

understand?

OpenStudy (anonymous):

see here u will hv to make the coefficient of either x ort y equal by multiplying by suitable nos and then subtracting second fr first as follows first by 3 : 27x + 6y = 6 (A) second by 1 : 21x + 6y = 4 (B) (A) - (B) gives 6x = 2 so x = 1/3 Put this in first eqn 9(1/3) + 2y = 2 3 + 2y = 2 2y = 2-3 y = -1/2

OpenStudy (anonymous):

ooh ok i get it thanks! now i can do the rest of the problems.

OpenStudy (angela210793):

U w :)